OpenStudy (dawnr):

HELP! what is the sum of digits in the number 13 in positional system with base 2?

OpenStudy (anonymous):

your picture is rad

OpenStudy (dawnr):

>_< can you help though?

OpenStudy (loser66):

let wait for @dan815 hihihi

OpenStudy (dawnr):

i just don't know what positional system is...i thought it was one of those that are done with an and sn but i don't know what to do with n and d...

OpenStudy (dawnr):

reffering to an=a1+(n-1)d and sn=n/2[2a1+(n-1)d]

OpenStudy (loser66):

ok, let me "interpret" the code for you. \(2^0 =1\\2^1=2\\2^2=4\\2^3=8\) and your number is 13, that means you have \(\color{red}{1}*2^3 +\color{red}{1}*2^2+\color{red}{0}*2^1+\color{red}{1}*2^0\), right? so, just pick the number in the front of 2^..., you have 1101 =13

OpenStudy (loser66):

dan is lazy guy. He is there but he doesn't want to do the easy stuff. :)

OpenStudy (loser66):

I don't get what you mean. \(13_{10}=1101_{2}\) if you want to find the sum of digits in base 2, it must be 3

OpenStudy (dawnr):

nvm the last comment..so basically i just need to guess the number..there isn't any formula or something like that..i just need to go from a number to a number

OpenStudy (loser66):

You don't understand the concept?? right??

OpenStudy (dawnr):

i do..but isn't there any other way..what if i have the number 50...or 146

OpenStudy (dawnr):

it's pain in the retriceto go step by step xD

OpenStudy (dawnr):

retrice*

OpenStudy (loser66):

the same!! just go from the last digit is 2^0, next is 2^1 .....

OpenStudy (dawnr):

do i always go to 2^3 or i can go to 2^n ?

OpenStudy (loser66):

sure.

OpenStudy (loser66):



OpenStudy (dawnr):

okay so i just need to find the number which is closest to the number i'm given

OpenStudy (loser66):

you have formula to find it out . This is "interpretation" the method, only. You understand what it is, then apply formula. In case you forget the formula, draw the table out and "defeat" the problem. The same for base 3,4,5...... and so on. Just remember that base2 uses 2 digits 0,1 base3 uses 3 digits 0,1.2 and so on

OpenStudy (dawnr):

right,i will!i did this such a long time ago and i'm just starting to remember things :D

OpenStudy (loser66):

I have to confess that I don't remember the formula, either.

OpenStudy (dawnr):

i tried to find it on the internet but i wasn't so lucky..and i'm searching it through my notebooks but it's pretty much the same thing

OpenStudy (loser66):

by that interpretation, I am ok. no need to remember too much. hihihii

OpenStudy (anonymous):

idk mane

OpenStudy (dawnr):

well i kinda need to remember everything since i have this big sort of test for collage in 2 weeks...and i haven't went past everything..so i'm just trying to catch up xD

OpenStudy (loser66):

http://en.wikipedia.org/wiki/Binary_number

OpenStudy (dawnr):

'cause i knew these things but totally forgot about them >_< *pelletty professor syndrome *

OpenStudy (dawnr):

s h i t t y*